OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is... y - 4 = 2/3 (x - 3) Anyone know how to solve this??

OpenStudy (bibby):

what do you know about the sloeps of parallel lines?

OpenStudy (bibby):

slopes*

OpenStudy (anonymous):

aren't they the same?

OpenStudy (bibby):

correct. and so the slope of our line is the same slope as the given line

OpenStudy (anonymous):

So, it is 2/3 ?

OpenStudy (bibby):

correct. now since we'd have to use point-slope form, we need one other piece of information \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

I'm not sure what it is.

OpenStudy (bibby):

(x1,y1) is the point the line goes through m is the slope of the line

OpenStudy (bibby):

we found m

OpenStudy (anonymous):

oh, so (1,-2) ?

OpenStudy (bibby):

correct. \(y-y_1=\dfrac{2}{3}(x-x_1)\)

OpenStudy (anonymous):

so, the equation would look like this? -2-4= 2/3 (1-3)

OpenStudy (bibby):

oh no, we're only replacing x1 and y1

OpenStudy (bibby):

the only information we needed from the other equation was the slope

OpenStudy (bibby):

\((x_1,y_1) = Q( 1, -2) \) m=2/3

OpenStudy (anonymous):

okay, i'm lost again. haha so how do we solve that?

OpenStudy (bibby):

lemme try some color coding

OpenStudy (bibby):

\(y-\color{red}{y_1}=\color{pink}m(x-\color{blue}{x_1})\) \(y - 4 = \color{pink}{\dfrac{2}{3}}(x - 3)\) passes through point \(Q( \color{blue}{1}, \color{red}{-2}) \)

OpenStudy (anonymous):

It makes some sense but i'm not sure how to solve it now

OpenStudy (bibby):

you just plug and simplify into y=mx+b form pretend it's one of those kindergarten toys where the red shape goes in the red hole

OpenStudy (anonymous):

y- (-2)= 2/3 (x-1) ?

OpenStudy (bibby):

yeah. start with the left hand side simplify y--2

OpenStudy (anonymous):

I got y= 2/3x-8/3 when I put it into standard form I got 2x-3y= 8 Which fits into my answer box online

OpenStudy (bibby):

why'd you put it in standard form?

OpenStudy (anonymous):

that's what the question said to do

OpenStudy (bibby):

oh I didn't read the first line, give me a moment

OpenStudy (anonymous):

it said I got it right so thank you! you helped me pass (:

OpenStudy (bibby):

y- (-2)= 2/3 (x-1) ? y+2=2/3x-2/3, subtract 6/3 (2) from both sides y=2/3x-8/3

OpenStudy (bibby):

awesome

OpenStudy (bibby):

i just googled how to convert to standard form for completion's sake y=2/3x-8/3 3y=3(2/3x-8/3) 3y=2x-8 2x-3y=8 good job

OpenStudy (anonymous):

Thank you! and thanks for the help also

OpenStudy (bibby):

anytime, yo. tag me in the future if it takes a while to get an answer